I need help bad with this!
Midtown Delivery Service delivers packages with cost $1.30 per package to deliver. The fixed cost to run the delivery truck is $75 per day. If the company charges $6.30 per package, how many packages must be delivered daily to break even?

Answer by rfer(16322)

1.3x+75=6.3x
5x=75
x=15
----------
1.3(15)+75=6.3(15)
94.5=94.5